OpenStudy (shakeria98):

What is the y-value of the solution to the system of equations? 3x + 5y = 1 7x + 4y = −13

OpenStudy (wanell):

3x + 5y = 1 /*(-4) 7x + 4y = - 13 /*5 ______________ -12x - 20y = -4 35x + 20y = -65 ______________ -12x - 20y + 35x + 20y = - 4 - 65 35x - 12x = - 69 23x = - 69 x = - 69/23 x = - 3 3x + 5y = 1 3 * (-3) + 5y = 1 -9 + 5y = 1 5y = 1 + 9 5y = 10 /5 y = 10/5 y = 2 The correct result would be y = 2.

OpenStudy (will.h):

2 ways to solve Elimination method Substitution method Which one you prefer